# BNSR-Invariants of Surface Houghton Groups

### v1

@@ 1 Introduction
To any group $G$, one can assign a sequence of invariants [^1] Technically, $\Sigma^{k}$ is only defined for groups of type $F_{k}$, but one could simply interpret this as empty when $G$ faisl to be of type $F_{k}$. $\Sigma^{k}(G)$ These invariants determine which subgroups of $G$ (containing the commutator) share which finiteness properties of $G$. Historically, they are difficult to compute. They were defined across several papers (see [BNS87], [BR88], and [Ren88]), primarily by Bieri, Neumann, Strebel, and Renz, hence the names “BNS-invariants” and “BNSR-invariants”. For the remainder of this paper, these will be referred to as $\Sigma$-invariants for the sake of brevity. One recent collection of groups for which these have been computed is the family of Houghton groups, in [Zar16] and [Zar20].
Houghton defined his groups as permutation groups of infinite sets in [Hou78]. More specifically, $H_{n}$ is the group of “eventual translations” of the set $\{1,\dots,n\}\times\mathbb{N}$, i.e. permutations which in each ray are translations outside some finite set. Brown proved, via a suitable simplicial complex, that Houghton’s group $H_{n}$ is of type $F_{n-1}$, but not of type $F_{n}$ (in the language of “finiteness length” this is $fl(H_{n})=n-1$) [^2] Actually, Brown proves type $FP_{n-1}$ but not type $FP_{n}$, and finitely presented, which together imply $F_{n}$ but not $F_{n-1}$.. In [ABKL23], the authors define a variant of Houghton groups, called surface Houghton groups, as asymptotically rigid mapping class groups of surfaces with infinite genus. We denote the surface Houghton groups by $\mathcal{H}_{n}$, and the pure (i.e. end-fixing) subgroups as $P\mathcal{H}_{n}$. In the same paper, they also show that $fl(\mathcal{H}_{n})=n-1$, via a cube complex analogous to the Stein–Farley complex.
@@
Let $\chi$ be a non-zero character of $P\mathcal{H}_{n}$. Then $[\chi]\in\Sigma^{m(\chi)-1}(P\mathcal{H}_{n})\setminus\Sigma^{m(\chi)}(P\mathcal{H}_{n})$.
As an application of this, we provide a partial converse (for Houghton groups and Surface Houghton groups) of the well-known fact that if $H\leq G$ is a finite index subgroup, then $fl(H)=fl(G)$. Specifically, we show that whenever a subgroup of $H_{n}$ or $P\mathcal{H}_{n}$ has finiteness length $n-1$ and intersects the corresponding commutator with finite index, it is finite index in the full group. Further, in order to be finite index, the intersection condition must hold, for general group theoretic reasons. (These are also true of $\mathcal{H}_{n}$, but for trivial reasons: the commutator of $\mathcal{H}_{n}$ is all of $\mathcal{H}_{n}$.) This is carried out in Section 5, along with some discussion of co-Hopfianness.
@@
To do this, we use a refinement of a proposition from [GLU20], see Proposition 3.2. In Section 2, we lay out various definitions and lemmas that we need, of the surfaces and groups we are concerned with, of the corresponding cube complex, and of the version of discrete Morse theory we shall employ. We also obtain various nice representatives for the vertices (Section 2) and edges (Section 3) of the complex.
@@
A recent paper by Marie Abadie [Aba24] analyses the CAT(0) property for the Stein–Farley cube-complexes of a different family of asymptotically rigid mapping class groups, including the braided Houghton groups. Similar methods are used, in particular a version of the same proposition as we use. Thus, one could attempt to apply Zaremsky’s methods to compute the $\Sigma$-invariants of $\operatorname{Br}H_{n}$ as well. It seems reasonable to guess that they should work out the same.
@@ Acknowledgements
We would like to thank Jing Tao for advice and guidance on this project. We would also like to thank Noel Brady and Justin Malestein for valuable conversations, and Peter Patzt and Matt Zaremsky for helpful comments. The authors were partially supported by the OU Bridge Funding Investment Program.
@@ 2.1 The (Pure) Surface Houghton Group
Here we lay out the definitions necessary for the group $P\mathcal{H}_{n}$, the pure version of the Surface Houghton Group defined in [ABKL23]. Let $\mathcal{O}=\mathcal{O}_{n}$ be a sphere with $n$ boundary components, and let $T$ be a torus with two boundary components, $\partial^{-}$ and $\partial^{+}$. Fix an orientation-reversing homeomorphism $\lambda:\partial^{-}\rightarrow\partial^{+}$, and for each $i$ an orientation-reversing homeomorphism $\mu_{i}$ from $\partial^{-}$ to the $i$th boundary components of $\mathcal{O}$. We construct $\Sigma_{n}$ as follows: begin with $M^{1}=\mathcal{O}$, then glue a copy of $T$ to each boundary component of $M^{1}$ via the $\mu_{i}$ to obtain $M^{2}$. For each $j\geq 2$, glue a copy of $T$ to each boundary component of $M^{j}$ via $\lambda$ to obtain $M^{j+1}$. Then the surface $\Sigma_{n}$ is the union of all the $M^{j}$’s; $\Sigma_{n}$ is the surface with $n$ ends, all accumulated by genus. Each of the closures of the components of $M^{j}\setminus M^{j-1}$ is called a piece, and each piece $B$ has a canonical homeomorphism $\iota_{B}:B\rightarrow T$. We will occasionally write $B_{k}^{j}$ to denote the $j$th piece in the $k$th end.
@@
We define some special elements of $\mathcal{H}_{n}$. For $\sigma$ a permutation of $\{1,\dots,n\}$, choose some homeomorphism of $\mathcal{O}$ permuting the boundary components in the same fashion. For the sake of definiteness, consider $\sigma$ first as an element of $\operatorname{Map}(S_{0,n})$, the sphere with $n$ punctures; then $\sigma$ yields a homeomorphism of $\mathcal{O}$, which we also dentoe by $\sigma$. Up to some isotopy, we can assume that the restriction of $\sigma$ to the $i$th boundary component of $\mathcal{O}$ is $\mu_{\sigma(i)}\circ\mu_{i}^{-1}$. Thus, we can extend this homeomorphism to be rigid outside of $\mathcal{O}$; call this extension again by the same name, $\sigma$. Note that such a map is not unique: there are many choices of mapping class in the sphere with $n$ punctures which induce the same permutation.
Next, we define the handle shifts $\rho_{i}$, $\grave{a}$ $la$ [PV18]. For $i\in\{1,\dots,n-1\}$, we want $\rho_{i}$ to be a homeomorphism of $\Sigma_{n}$ which shifts the $i$th end towards the $n$th end by a single genus. Specifically, map each piece $B_{n}^{j}$ to $B_{n}^{j+1}$, and for $j\geq 1$ map each piece $B_{i}^{j+1}$ to $B_{i}^{j}$; choose these to be the rigid maps. The ends other than $i$ and $n$ are unchanged, so all that remains is to define how $\rho_{i}$ acts on $B_{i}^{1}\cup\mathcal{O}$. This is demonstrated by figure 2.1 Note that $\rho_{i}$ is asymptotically rigid outside of $\mathcal{O}\cup B_{i}^{1}$.
{tikzpicture}
Figure 1: A handle shift being applied to a select curve.
@@
Consider an arbitrary element $\varphi\in\mathcal{H}_{n}$: $\varphi$ clearly induces some permutation $\sigma$ on the ends of $\Sigma_{n}$; obtain from this $\sigma\in\mathcal{H}_{n}$. Now we have that $\varphi\circ\sigma^{-1}\in P\mathcal{H}_{n}$, so we wish to analyse how pure mapping classes act within ends. A rigorous examination of this is contained in Section 2.3. In the meantime, consider how $\varphi$ acts in the complement of a defining surface: as it does not permute the ends, and takes pieces to pieces, it must shift each end in or out by some integer amount. From this, we obtain a sequence of handle shifts $\rho_{i_{1}},\dots,\rho_{i_{k}}$. Finally, we have that $\varphi\circ\sigma^{-1}\circ(\rho_{i_{1}}\cdots\rho_{i_{k}})^{-1}$ is a compactly supported mapping class, $\alpha$. ∎
@@
The above expression obtained for $\varphi$ is far from unique: choices were made both in the selection of $\sigma$, as well as in the selection of the handle shifts. However, any different choices made for these will simply change $\alpha$, and the properties of $\alpha$ are rarely (never?) relevant for this paper.
@@ 2.2 The Contractible Cube Complex
Denote the equivalence class of $(Z,\varphi)$ by $[Z,\varphi]$, and the set of equivalence classes by $SS\SS$. The group $\mathcal{H}_{n}$ acts on $SS\SS$ by
@@
Define the complexity of a pair $(Z,\varphi)$ to be the genus of $Z$. This extends to be a $\mathcal{H}_{n}$-invariant height function $SSf\colon\SS\rightarrow\mathbb{Z}\subset\mathbb{R}$. Given vertices $SSx_{1},x_{2}\in\SS$, we say that $x_{1}\prec x_{2}$ if there are representatives $(Z_{i},\varphi_{i})$ of $x_{i}$ so that $\varphi_{1}=\varphi_{2}$, $Z_{1}\subset Z_{2}$, and $\overline{Z_{2}\setminus Z_{1}}$ is a disjoint union of pieces.
We now construct $X_{n}$ as a cube complex with vertex set $SS\SS$. We declare that there is a $d$-cube whenever $x_{1}\prec x_{2}$, with $d=f(x_{2})-f(x_{1})$. Note that as $\Sigma_{n}$ has $n$ ends, the complex $X_{n}$ is $n$ dimensional. The height function $f$ extends affinely to a $\mathcal{H}_{n}$-invariant height function on $X_{n}$. Also, let $X_{n}^{f\leq k}$ denote the subcomplex spanned by vertices with height at most $k$. We have the following properties (see [ABKL23, Theorem 4.1]):
@@
Any edge in $X_{n}$ can be expressed as $[Z,\varphi]$$[Z\cup B,\varphi]$, where $\varphi\in P\mathcal{H}_{n}$.
@@
Consider a vertex $[Z,\varphi]$, and let $B$ be a piece adjacent to $Z$, so that $Z\cup B$ is suited. Write $\varphi=\alpha\rho_{i_{1}}\cdots\rho_{i_{j}}\sigma$ as in Lemma 2.2. Then we wish to show that $[Z,\varphi]=[\sigma(Z),\varphi\circ\sigma^{-1}]$, and that $[Z\cup B,\varphi]=[\sigma(Z)\cup\sigma(B),\varphi\circ\sigma^{-1}]$. Consider the following diagram.
{tikzpicture}
As $\sigma$ is rigid away from the center, we have that $\sigma(B)$ is a piece, and of course it must be adjacent to $\sigma(Z)$, so all that remains to be checked is that $[Z,\varphi]=[\sigma(Z),\varphi\circ\sigma^{-1}]$. First, note that $(\varphi\circ\sigma^{-1})^{-1}\circ\varphi=\sigma$, which is rigid away from $\mathcal{O}$, and thus takes $Z$ to $\sigma(Z)$ (a suited subsurface), and is certainly rigid outside of $Z$. ∎
@@
Any edge in $X_{n}$ can be expressed as $[L_{g},\varphi]$$[L_{g+1},\psi]$.
@@
Suppose we are given a vertex $[Z,\varphi^{\prime}]$. Recall that the handle shift $\rho_{i}$ is rigid away from $B_{i}^{1}$, the piece that it pushes across the center. As $(\varphi)^{-1}\circ\varphi^{\prime}$ can be chosen to be a composition of handle shifts taking $Z$ to $L_{g}$, the only places where $(\varphi)^{-1}\circ\varphi^{\prime}$ can be non-rigid are pieces in $Z$. This shows that vertices can be represented as $[L_{g},\varphi]$. To extend this to the edge $[Z,\varphi^{\prime}]$$[Z\cup B,\varphi^{\prime}]$, consider the following diagram.
{tikzpicture}
Observe that to obtain an $L_{g+1}$ form for $[Z\cup B,\varphi^{\prime}]$, we can first push $Z$ to $L_{g}$, which sends $B$ to a piece $B_{i}^{1}$, and then append one more $\rho_{i}$. ∎

### v2

@@ 1 Introduction
To any group $G$ of type $F_{k}$, one can assign a sequence of invariants $\Sigma^{m}(G)$, for $m\leq k$. These invariants determine which subgroups of $G$ (containing the commutator subgroup) share which finiteness properties of $G$. Historically, they are difficult to compute. They were defined across several papers (see [BNS87], [BR88], and [Ren88]), primarily by Bieri, Neumann, Strebel, and Renz, hence the names “BNS-invariants” and “BNSR-invariants”. For the remainder of this paper, these will be referred to as $\Sigma$-invariants for the sake of brevity. One recent collection of groups for which these have been computed is the family of Houghton groups, in [Zar16] and [Zar20].
Houghton defined his groups as permutation groups of infinite sets in [Hou78]. More specifically, $H_{n}$ is the group of “eventual translations” of the set $\{1,\dots,n\}\times\mathbb{N}$, i.e. permutations which in each ray are translations outside some finite set. Brown proved, via a suitable simplicial complex, that Houghton’s group $H_{n}$ is of type $F_{n-1}$, but not of type $F_{n}$ (in the language of “finiteness length” this is $fl(H_{n})=n-1$) [^1] Brown proves type $FP_{n-1}$ but not type $FP_{n}$, and finitely presented, which together imply $F_{n}$ but not $F_{n-1}$.. In [ABKL23], the authors define a variant of Houghton groups, called surface Houghton groups, as asymptotically rigid mapping class groups of surfaces with infinite genus. We denote the surface Houghton groups by $\mathcal{H}_{n}$, and the pure (i.e. end-fixing) subgroups as $P\mathcal{H}_{n}$. In the same paper, they also show that $fl(\mathcal{H}_{n})=n-1$, via a cube complex analogous to the Stein–Farley complexes for Thompson groups.
@@
Let $\chi$ be a non-zero character of $P\mathcal{H}_{n}$. Then $\chi]\in\Sigma^{m(\chi)-1}(P\mathcal{H}_{n})\setminus\Sigma^{m(\chi)}(P\mathcal{H}_{n})$.
As an application of this, we provide a partial converse (for Houghton groups and surface Houghton groups) of the well-known fact that if $H\leq G$ is a finite index subgroup, then $fl(H)=fl(G)$. Specifically, we show that whenever a subgroup of $H_{n}$ or $P\mathcal{H}_{n}$ has finiteness length n1n-1 and intersects the corresponding commutator with finite index, it is finite index in the full group. Further, in order to be finite index, the intersection condition must hold, for general group theoretic reasons. (These are also true of $\mathcal{H}_{n}$, but for trivial reasons: the commutator subgroup of $\mathcal{H}_{n}$ is all of $\mathcal{H}_{n}$.) This is carried out in Section 5, along with some discussion of co-Hopfianness.
@@
To do this, we use a refinement of a proposition from [GLU20], see Proposition 3.2. In Section 2, we lay out the definitions of the surfaces and groups we are concerned with, of the corresponding cube complex, and of the version of discrete Morse theory we shall employ. Alongside these definitions are various lemmas which we shall need. We also obtain various nice representatives for the vertices (Section 2) and edges (Section 3) of the complex.
@@
A recent paper by Marie Abadie [Aba24] analyzes the CAT(0) property for the Stein–Farley cube complexes of a different family of asymptotically rigid mapping class groups, including the braided Houghton groups. Similar methods are used, in particular a version of the same proposition as we use. Thus, one could attempt to apply Zaremsky’s methods to compute the $\Sigma$-invariants of $\operatorname{Br}H_{n}$ as well. It seems reasonable to guess that they should work out the same.
@@ Acknowledgements
We would like to thank Jing Tao for advice and guidance on this project. We would also like to thank Noel Brady and Justin Malestein for valuable conversations, and Peter Patzt and Matt Zaremsky for helpful comments. We also thank the anonymous referee for helpful comments. The authors were partially supported by the OU Bridge Funding Investment Program.
@@ 2.1 The (Pure) Surface Houghton Group
Here we lay out the definitions necessary for the group $P\mathcal{H}_{n}$, the pure version of the surface Houghton group defined in [ABKL23]. Let $\mathcal{O}=\mathcal{O}_{n}$ be a sphere with $n$ boundary components, and let $T$ be a torus with two boundary components, $\partial^{-}$ and $\partial^{+}$. Fix an orientation-reversing homeomorphism $\lambda:\partial^{-}\rightarrow\partial^{+}$, and for each $i$ an orientation-reversing homeomorphism $\mu_{i}$ from $\partial^{-}$ to the $i$th boundary component of $\mathcal{O}$. We construct $\Sigma_{n}$ as follows: begin with $M^{1}=\mathcal{O}$, then glue a copy of $T$ to each boundary component of $M^{1}$ via the $\mu_{i}$ to obtain $M^{2}$. For each $j\geq 2$, glue a copy of $T$ to each boundary component of $M^{j}$ via $\lambda$ to obtain $M^{j+1}$. Then the surface $\Sigma_{n}$ is the union of all the $M^{j}$’s; $\Sigma_{n}$ is the surface with $n$ ends, all accumulated by genus. We call $\mathcal{O}$ the center of $\Sigma_{n}$, each of the closures of the components of $M^{j}\setminus M^{j-1}$ is called a piece, and each piece $B$ has a canonical homeomorphism $\iota_{B}:B\rightarrow T$. We will occasionally write $B_{k}^{j}$ to denote the $j$th piece in the $k$th end.
@@
We define some special elements of $\mathcal{H}_{n}$. For $\sigma$ a permutation of $\{1,\dots,n\}$, choose some homeomorphism of $\mathcal{O}$ permuting the boundary components in the same fashion. For the sake of definiteness, consider $\sigma$ first as an element of $\operatorname{Map}(S_{0,n})$, the sphere with $n$ punctures; then $\sigma$ yields a homeomorphism of $\mathcal{O}$, which we also denote by $\sigma$. Up to some isotopy, we can assume that the restriction of $\sigma$ to the $i$th boundary component of $\mathcal{O}$ is $\mu_{\sigma(i)}\circ\mu_{i}^{-1}$. Thus, we can obtain a homeomoprhism of $\Sigma_{n}$ by extending this map to be rigid outside of $\mathcal{O}$; call this extension again by the same name, $\sigma$. Note that such a map is not unique: there are many choices of mapping class in the sphere with $n$ punctures which induce the same permutation.
Next, we define the handle shifts $\rho_{i}$, $\grave{a}$ lala [PV18]. For $i\in\{1,\dots,n-1\}$, we want $\rho_{i}$ to be a homeomorphism of $\Sigma_{n}$ which shifts the $i$th end towards the $n$th end by a single genus. Specifically, for $j\geq 1$ map each piece $B_{n}^{j}$ to $B_{n}^{j+1}$, and map each piece $B_{i}^{j+1}$ to $B_{i}^{j}$; choose these to be the rigid maps. The ends other than $i$ and $n$ are unchanged, so all that remains is to define how $\rho_{i}$ acts on $B_{i}^{1}\cup\mathcal{O}$. This is demonstrated by Figure 1. Note that $\rho_{i}$ is asymptotically rigid outside of $\mathcal{O}\cup B_{i}^{1}$.
$\overset{\rho_{i}}{\longrightarrow}iijjnn$
Figure 1: The handle shift $\rho_{i}$ (which pushes from end $i$ to end $n$) being applied to various curves in ends $i$, $n$, and $j$.
@@
Consider an arbitrary element $\varphi\in\mathcal{H}_{n}$: $\varphi$ clearly induces some permutation $\sigma$ on the ends of $\Sigma_{n}$; obtain from this $\sigma\in\mathcal{H}_{n}$. Now we have that $\varphi\circ\sigma^{-1}\in P\mathcal{H}_{n}$, so we wish to analyze how pure mapping classes act within ends. A rigorous examination of this is contained in Section 2.3. In the meantime, consider how $\varphi$ acts in the complement of a defining surface: as it does not permute the ends, and takes pieces to pieces, it must shift each end in or out by some integer amount. From this, we obtain a sequence of handle shifts $\rho_{i_{1}},\dots,\rho_{i_{k}}$. Finally, we have that $\varphi\circ\sigma^{-1}\circ(\rho_{i_{1}}\cdots\rho_{i_{k}})^{-1}$ is a compactly supported mapping class, $\alpha$. ∎
@@
The above expression obtained for $\varphi$ is far from unique: choices were made both in the selection of $\sigma$, as well as in the selection of the handle shifts. However, any different choices made for these will simply change $\alpha$, and the properties of $\alpha$ are rarely relevant for this paper.
@@ 2.2 The Contractible Cube Complex
Denote the equivalence class of $(Z,\varphi)$ by $Z,\varphi]$, and the set of equivalence classes by $S\SS$. The group $\mathcal{H}_{n}$ acts on $S\SS$ by
@@
Define the complexity of a pair $(Z,\varphi)$ to be the genus of $Z$. This extends to be an $\mathcal{H}_{n}$-invariant height function $Sf\colon\SS\rightarrow\mathbb{Z}\subset\mathbb{R}$. Given vertices $Sx_{1},x_{2}\in\SS$, we say that $x_{1}\prec x_{2}$ if there are representatives $(Z_{i},\varphi_{i})$ of $x_{i}$ so that $\varphi_{1}=\varphi_{2}$, $Z_{1}\subset Z_{2}$, and $\overline{Z_{2}\setminus Z_{1}}$ is a (non-empty) disjoint union of pieces (i.e. has at most one piece from each end).
We now construct $X_{n}$ as a cube complex with vertex set $S\SS$. We declare that whenever $x_{1}\prec x_{2}$, with $d=f(x_{2})-f(x_{1})$, there is a $d$-cube with vertex set given by $\{x\,|\,x_{1}\preceq x\preceq x_{2}\}$. Note that as $\Sigma_{n}$ has $n$ ends, the complex $X_{n}$ is $n$ dimensional. The height function $f$ extends affinely to an $\mathcal{H}_{n}$-invariant height function on $X_{n}$. Also, let $X_{n}^{f\leq k}$ denote the subcomplex spanned by vertices with height at most $k$. We have the following properties (see [ABKL23, Theorem 4.1]):
@@
Any edge in $X_{n}$ can be expressed as $Z,\varphi][$$Z\cup B,\varphi]$, where $\varphi\in P\mathcal{H}_{n}$.
@@
Consider a vertex $Z,\varphi]$, and let $B$ be a piece adjacent to $Z$, so that $Z\cup B$ is suited. Write $\varphi=\alpha\rho_{i_{1}}\cdots\rho_{i_{j}}\sigma$ as in Lemma 2.2. Then we wish to show that $Z,\varphi]=[\sigma(Z),\varphi\circ\sigma^{-1}]$, and that $Z\cup B,\varphi]=[\sigma(Z)\cup\sigma(B),\varphi\circ\sigma^{-1}]$. Consider the following diagram.
$\sigma(Z),\varphi\circ\sigma^{-1}][$$\sigma(Z\cup B),\varphi\circ\sigma^{-1}][$$\sigma(Z)\cup\sigma(B),\varphi\circ\sigma^{-1}][$$Z,\varphi][$$Z\cup B,\varphi]$
As $\sigma$ is rigid away from the center $\mathcal{O}$, we have that $\sigma(B)$ is a piece, and of course it must be adjacent to $\sigma(Z)$, so all that remains to be checked is that $Z,\varphi]=[\sigma(Z),\varphi\circ\sigma^{-1}]$. First, note that $(\varphi\circ\sigma^{-1})^{-1}\circ\varphi=\sigma$, which is rigid away from $\mathcal{O}$, and thus takes $Z$ to $\sigma(Z)$ (a suited subsurface), and is certainly rigid outside of $Z$. ∎
@@
Any edge in $X_{n}$ can be expressed as $L_{g},\varphi][$$L_{g+1},\psi]$.
@@
Suppose we are given a vertex $Z,\varphi^{\prime}]$. Recall that the handle shift $\rho_{i}$ is rigid away from $B_{i}^{1}$, the piece that it pushes across the center. As $(\varphi)^{-1}\circ\varphi^{\prime}$ can be chosen to be a composition of handle shifts taking $Z$ to $L_{g}$, the only places where $(\varphi)^{-1}\circ\varphi^{\prime}$ can be non-rigid are pieces in $Z$. This shows that vertices can be represented as $L_{g},\varphi]$. To extend this to the edge $Z,\varphi^{\prime}][$$Z\cup B,\varphi^{\prime}]$, consider the following diagram.
$L_{g},\varphi][$$L_{g}\cup B_{i}^{1},\varphi][$$L_{g+1},\varphi\circ\rho_{i}^{-1}]$
Observe that to obtain an $L_{g+1}$ form for $Z\cup B,\varphi^{\prime}]$, we can first push $Z$ to $L_{g}$, which sends $B$ to a piece $B_{i}^{1}$, and then append one more $\rho_{i}$. ∎